OpenStudy (anonymous):

Factor completely. x 3 + x 2 - 25x - 25 (x+ 1)(x- 5)(x+ 5) (x+ 5)(x- 1)(x+ 5) (x- 1)(x+ 5)(x- 5)

OpenStudy (anonymous):

@t0mmiere

OpenStudy (anonymous):

You jumped from grade 7 to grade 12 haha.. (x+5)(x-5)(x+1)

OpenStudy (anonymous):

The answer here is correct. The method used is factoring by grouping.\[x^3 + x^2 - 25x - 25=x^2(x+1)-25(x+1)=(x^2-25)(x+1)\]\[=(x+1)(x-5)(x+5)\]